OpenStudy (anonymous):

Please help! Will award medal! Find an equation for the nth term of the arithmetic sequence. a19 = -92, a20 = 6

OpenStudy (anonymous):

@Abhisar

OpenStudy (asnaseer):

do you know the formula for the n'th term of an arithmetic sequence?

OpenStudy (asnaseer):

alternatively, do you know what an arithmetic sequence is?

OpenStudy (anonymous):

no and how would i use the 19th and 20th term anyway? @asnaseer

OpenStudy (asnaseer):

what is the property of an arithmetic sequence - what "makes it" an arithmetic sequence?

OpenStudy (anonymous):

the pattern?

OpenStudy (anonymous):

so adding 98?

OpenStudy (asnaseer):

correct - so an arithmetic sequence is a sequence you get by adding the same number to every term (the number you add could be negative)

OpenStudy (asnaseer):

in your case you are given the 19th and 20th terms, so you can say that:\[a_{20}=a_{19}+d\]where 'd' is the difference between each term in the sequence - make sense?

OpenStudy (anonymous):

so 98?

OpenStudy (asnaseer):

correct - so now you know that this "common difference" is 98 You now need to recall the formula for the n'th term

OpenStudy (anonymous):

ok so its -1856 and 98(n-1)

OpenStudy (anonymous):

but is it + or - the 98(n-1)

OpenStudy (asnaseer):

imagine you had such a sequence:\[a_{1},a_{2},a_{3},...\]and you know that the common difference between them is 'd'. then you could rewrite this sequence as:\[a_1,a_1+d,a_1+2d,...\]agreed?

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

if you compare this to the original sequence of: \(a_1,a_2,a_3,...\) you will notice that we can say that:\[a_n=a_1+(n-1)d\]this gives us the formula for the n'th term in the sequence

OpenStudy (asnaseer):

we already know what 'd' is but we do not yet know what \(a_1\) is equal to

OpenStudy (asnaseer):

so what you can do is use the information given to you - e.g. you are told that the 20th term is 6, therefore we can write:\[a_{20}=6=a_1+(20-1)d=a_1+(20-1)\times98\]

OpenStudy (asnaseer):

use this to work out what \(a_1\) is equal to

OpenStudy (anonymous):

isnt it -1856

OpenStudy (asnaseer):

perfect!

OpenStudy (asnaseer):

now put it all together and we get:\[a_n=a_1+(n-1)d=-1856+(n-1)98\]

OpenStudy (asnaseer):

as a sanity check you could use this to see if \(a_{19}\) and \(a_{20}\) come out correctly

OpenStudy (anonymous):

so -1856+98(n-1)

OpenStudy (asnaseer):

yes

OpenStudy (anonymous):

cool thanks!

OpenStudy (asnaseer):

yw :)

OpenStudy (anonymous):

i actually got it way earlier up there^ but you ended up explaining it cuz i was just guessing but thanks!

OpenStudy (anonymous):

do you mind helping me with another one?

OpenStudy (asnaseer):

it's always better to understand the concepts - it will make solving the next one much easier :)

OpenStudy (asnaseer):

I need to go and eat now - but just post a new question and I'm sure there will be plenty of good people here to help you

OpenStudy (anonymous):

haha okay thanks!